# Contextual Stochastic Optimization for Order Fulfillment Optimization

### v1

@@
Problem definition: The paper studies a large-scale order fulfillment problem for a leading e-commerce company in the United States. The challenge involves selecting fulfillment centers and shipping carriers with observational data only to efficiently process orders from a vast network of physical stores and warehouses. The company’s current practice relies on heuristic rules that choose the cheapest fulfillment and shipping options for each unit, without considering opportunities for batching items or the reliability of carriers in meeting expected delivery dates. Methodology / results: The paper develops a data-driven Contextual Stochastic Optimization (CSO) framework that integrates distributional forecasts of delivery time deviations with stochastic and robust order fulfillment optimization models. The framework optimizes the selection of fulfillment centers and carriers, accounting for item consolidation and delivery time uncertainty. Validated on a real-world data set containing tens of thousands of products, each with hundreds of fulfillment options, the proposed CSO framework significantly enhances the accuracy of meeting customer-expected delivery dates compared to current practices. It provides a flexible balance between reducing fulfillment costs and managing delivery time deviation risks, emphasizing the importance of contextual information and distributional forecasts in order fulfillment. Managerial implications: This is the first paper that studies the omnichannel multi-courier order fulfillment problem with delivery time uncertainty through the lens of contextual optimization, fusing machine learning and optimization. The results provide actionable insights for e-commerce companies and online retailers to enhance service quality and customer satisfaction through efficient order fulfillment strategies that accounts for delivery time uncertainty. Adopting this framework can balance cost effectiveness with customer service, potentially leading to higher customer retention and profitability.
@@
logistics, order fulfillment, machine learning, contextual stochastic optimization, contextual robust optimization
@@ 1 Introduction
Most existing studies on order fulfillment optimization have not explicitly considered the uncertainty in delivery timeliness. However, Freedman (2019) highlight the strong desire of customers for accurate delivery time promises and faster service. The survey indicates that around 44% of online shoppers have abandoned their shopping carts because the items they wanted would not arrive on time, while 20% abandon orders with unclear delivery dates. Furthermore, 28% of shoppers are willing to pay for expensive, expedited shipping to receive products at their desired time. The statistics clearly show that customers prioritize service quality, sometimes even more than the product value, when making purchase decisions. Moreover, Salari et al. (2022) point out the importance of considering delivery time uncertainty in order fulfillment to better meet customer expectations and improve satisfaction. Their proposed promised delivery time policy demonstrates a 6.1% increase in sales volume in numerical experiments. Notably, overpromising delivery times can lead to higher return rates, lower repurchase rates, and dissatisfaction with delayed service. Additionally, customers can have mixed sentiments over underpromised delivery times, as early deliveries may not align with their availability. Therefore, missing the exact promised delivery date can negatively impact short-term and long-term sales, both implicitly and explicitly (Chan et al. 2018, Cui et al. 2020, Salari et al. 2022). Given the crucial role of delivery timeliness, the paper incorporates delivery time uncertainty into the omnichannel multi-courier order fulfillment optimization problem through a contextual stochastic optimization (CSO) framework.
One of the key challenges of Omnichannel Multi-Courier Order Fulfillment Optimization comes from the nature of historical data: only observational data is available, i.e., only the delivery time deviations of the selected options are available in the data set. In addition, in real time, these deviations are only observable once the decisions have been taken. The availability of observation data only raises fundamental challenges that are typically not addressed in prior work.
@@ 1.1 Contributions
The key novelty in this paper is a CSO framework for an omnichannel multi-courier order fulfillment optimization with observational data only. Existing studies on CSO typically assume the availability of a dataset $\mathcal{D}=\{(\mathbf{s}_{1},\mathbf{c}_{1}),(\mathbf{s}_{2},\mathbf{c}_{2}),\ldots\}$, pairing observed covariates with complete uncertain parameter vectors. However, real-world scenarios often involve high-dimensional uncertain parameters where individual data points capture only partial information, particularly when uncertainty measurement depends on the chosen solution, which is the case in the application considered in this paper.
@@
    To the best of the authors’ knowledge, this paper is the first study of the order fulfillment problem in the context of omnichannel and multi-courier logistics through the lens of data-driven contextual stochastic optimization.
@@
    A mixed-integer linear program (MILP) is developed for a novel order fulfillment optimization problem that jointly considers the locations of fulfillment centers and the selection of shipping carriers for online order fulfillment. This MILP accounts for the consolidation of items within each order.
@@
    Two classes of machine learning models are proposed to learn the delivery time deviations: probabilistic multi-class classification models and quantile regression models that provide distributional forecasts of delivery time deviations.
@@
    The proposed MILP and machine learning models are integrated into a generalized CSO framework. The framework integrates the conditional distributions of delivery time deviations with stochastic and robust optimization models via sample average approximation (SAA) and tractable reformulations.
  5. 5.
    The paper applies the CSO framework to a large-scale real-world instance, demonstrating its impact on balancing reduced fulfillment costs and increased on-time order deliveries. The results validate the practical effectiveness of contextual optimization in an e-commerce setting, bringing significant benefits compared to the baselines.
@@ 1.2 Structure of the Paper
The remaining sections of the paper are organized as follows. Section 2 reviews the related literature. Section 3 formalizes the generic CSO problem addressed in this paper and proposes SAA and robust optimization methods. The methods are applied to an unique order fulfillment problem in Section 4, which describes the MILP model and the machine learning models. Section 5 describes the case study, provides experimental results, and discusses actionable insights obtained from the experiments based on a real dataset. Section 6 concludes the paper.
@@ 2.1 Data-Driven Decision Making
This paper adds to the growing body of literature on data-driven decision-making, particularly at the intersection of machine learning and optimization, which has led to the development of CSO. Sadana et al. (2024) put forth a comprehensive survey on contextual optimization. Tutorials on various methodologies and applications of CSO models in logistics and operations management are available in the works of Mišić and Perakis (2020), Qi and Shen (2022), and Tian et al. (2023). Among the body of research on CSO, the most relevant to this paper are those that focus on learning conditional distributions of uncertain parameters and incorporating these forecasts into optimization models. Data-driven approaches typically build these forecasts from empirical residual errors (Deng and Sen 2022, Kannan et al. 2023, Perakis et al. 2023) or weighted empirical distributions based on proximity to the training data (Bertsimas and Kallus 2020, Notz and Pibernik 2022). However, these studies assume full observation of uncertain parameters for each data point, which differs from the context of this paper. Recent research has also explored integrating downstream optimization effects into ML model training. Elmachtoub and Grigas (2022) introduce the Smart Predict then Optimize (SPO) loss and its convex surrogate for point prediction models, with subsequent extensions to conditional distributional estimation (Qi et al. 2021, Kallus and Mao 2023). However, the SPO loss framework is not directly applicable to this paper’s context, as the dataset analyzed here is observational. In such datasets, outcomes are only partially observed based on decisions made, presenting unique challenges not addressed by these existing methods.
This paper thus contributes to the field of data-driven decision-making that handles observational data, a domain explored by prior research, e.g., in Bertsimas et al. (2019) and Jo et al. (2021) using optimal prescriptive trees for prescriptive analytics in personalized treatment. While these methods effectively learn policies mapping features to actions, they cannot be directly applied to this fulfillment problem considered in this paper due to the presence of instance-specific hard constraints on decisions. The proposed framework addresses these limitations, offering a novel approach to CSO problems with observational data and constrained decision spaces.
@@ 2.2 Order Fulfillment Problem
3 Problem and Methodology
This section describes a generic contextual stochastic optimization problem (denoted by ${\cal P}$) for fulfillment applications where only an observational dataset is available. It then outlines the overall approach taken to address the problem.
3.1 Problem Description
Problem ${\cal P}$ consists in fulfilling a set of incoming requests sequentially, each of which contains orders for a set of products. Each request $o$ has an associated quantity $q_{io}$ for some product $i\in\mathcal{I}$. These products can be fulfilled from a set $\mathcal{V}$ of selection options. When a request $o$ arrives and $o$ contains a subset of products in $\mathcal{I}$, the goal is to determine the quantity of product $i$ assigned to selection option $v$. The decision variables can thus be captured by a vector $\mathbf{z}_{o}=(z_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$, where each $z_{ivo}$ indicates the quantity of product $i$ assigned to selection option $v$ in request $o$. Each such assignment must satisfy operational constraints and the set of feasible assignments is denoted by $\mathcal{Z}_{o}$. For each product $i$, each selection option $v$ is associated with an uncertain cost $\tilde{c}_{ivo}\in\mathbb{R}$. The full vector of uncertain costs, governed by a probability distribution $\mathbb{P}$, is denoted by $\mathbf{\tilde{c}}_{o}=(\tilde{c}_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$. The objective function $g(\mathbf{z}_{o},\mathbf{\tilde{c}_{o}})$ of the problem depends on these uncertain costs and the decision variables. Problem ${\cal P}$ is thus a risk-neutral stochastic optimization which can be formulated as follows for request $o$:
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\mathbf{\tilde{c}}_{o}\sim\mathbb{P}}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})],$ (1)
where $\mathbb{E}_{\mathbf{\tilde{c}}_{o}\sim\mathbb{P}}[\cdot]$ represents the expectation of the objective function over the probability distribution $\mathbb{P}$.
3.2 Contextual Stochastic Optimization (CSO)
In Problem (1), the distribution $\mathbb{P}$ is not readily available. However, side information is typically available for product $i$ and selection option $v$. Let the vector of covariates for request be $\mathbf{s}_{ivo}$. In the context of an omnichannel order fulfillment problem, these covariates may represent the state of the fulfillment network, the weather, and the current demand for products. All these covariate vectors are aggregated into state vector $\mathbf{s}_{o}=(\mathbf{s}_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$, which encompasses all contextual information available to the request at the time of the assignment decision.
The distribution $\mathbb{P}$ is then approximated by the conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ of the cost vector given the observed state vector. This leads to the following Contextual Stochastic Optimization (CSO) problem:
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})],$ (2)
where $\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}[\cdot]$ is the conditional expectation of the objective function given the state vector $\mathbf{s}_{o}$. Let the uncertain cost vector and the aggregated covariates for selection option $v$ be $\tilde{\mathbf{c}}_{vo}=(\tilde{c}_{ivo})_{i\in\mathcal{I}}$ and $\mathbf{s}_{vo}=(\mathbf{s}_{ivo})_{i\in\mathcal{I}}$, respectively. This paper assumes that option costs are independent, i.e.,
$\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})=\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo}),$
where $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})$ is the conditional distribution of the uncertain cost $\tilde{\mathbf{c}}_{vo}$ given the covariates $\mathbf{s}_{vo}$. This assumption is reasonable as the uncertain cost for a specific option depends solely on the contextual features of that option, not on the contextual features of other options. Note that, although these conditional distributions are considered independent, selection options that share common covariates (e.g., weather) and thus their distributions will tend to be similar. Problem (2) can then be rewritten as
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})].$ (3)
3.3 Learning the Contextual Distribution
The conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ is also generally unknown but, in practice, it can often be learned from historical data available. In many real-world settings however, each data point is observational, i.e., only those costs involved in the actual decision can be observed. As a result, the dataset available to estimate $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ is of the form $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{C}_{o})\}_{o=1}^{|\mathbb{D}|}$, where $\mathbf{C}_{o}$ are the observable costs for order $o$. The learning task consists in finding a machine learning $\mathcal{M}$ that approximates $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ in a supervised manner using the dataset $\mathbb{D}$: $\mathcal{M}$ receives a context $\mathbf{s}$ as input and returns a probability distribution $\mathcal{M}(\mathbf{s})$ of costs. Since learning $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ amounts to learning $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})$ for each option $v$, the dataset $\mathbb{D}$ is partitioned into subsets, with $\mathbb{D}_{v}=\{(\mathbf{s}_{v1},\mathbf{c}_{v1}),(\mathbf{s}_{v2},\mathbf{c}_{v2}),\ldots\}$ representing the dataset for option $v$. A machine learning model $\mathcal{M}_{v}$ is associated with each option $v$ and $\mathcal{M}(\mathbf{\mathbf{s}})=(\mathcal{M}_{1}(\mathbf{s_{1}}),\ldots,\mathcal{M}_{|\mathcal{V}|}(\mathbf{s}_{|\mathcal{V}|}))$.
3.4 Contextual Sample Average Approximation (C-SAA)
Problem (3) is approximated by generalizing the Sample Average Approximation (SAA) method Kleywegt et al. (2002) to contextual stochastic optimization. The core of the Contextual Sample Average Approximation (C-SAA) for a request $o$ with covariates $\mathbf{s}_{o}$ amounts to
  1. 1.
    sampling the distribution $\mathcal{M}(\mathbf{s}_{o})$ $N$ times to obtain a set $\bm{\xi}_{o}^{n}=(\bm{\xi}_{vo}^{n})_{v\in\mathcal{V}}$ of cost vectors $(n\in[N])$;
  2. 2.
    solving the C-SAA problem
    (C-SAA) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\frac{1}{N}\sum_{n=1}^{N}g(\mathbf{z}_{o},\bm{\xi}_{o}^{n}).$ (4)
Following Kleywegt et al. (2002), the C-SAA procedure can be outlined as follows:
  1. 1.
    Solve $Q$ independent C-SAA problems, each of sample size $N=N_{1}$, to produce candidate solutions $\mathbf{z}_{1},\ldots,\mathbf{z}_{Q}$ with corresponding objective values $\underaccent{\bar}{f}_{1},\ldots\underaccent{\bar}{f}_{Q}$. Calculate the average objective value as $\underaccent{\bar}{f}=\frac{1}{Q}\sum_{i=1}^{Q}\underaccent{\bar}{f}_{i}$. Let $f^{*}$ be the true optimal objective value of Problem (2). Since $\underaccent{\bar}{f}$ is an unbiased estimator of $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}}[\underaccent{\bar}{f}]$, and because $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}}[\underaccent{\bar}{f}]\leq f^{*}$, $\underaccent{\bar}{f}$ serves as a statistical lower bound on $f^{*}$ (Kleywegt et al. 2002).
  2. 2.
    Then, re-evaluate the solutions $\mathbf{z}_{1},\ldots,\mathbf{z}_{Q}$ using a larger sample size $N=N_{2}\gg N_{1}$, yielding new objective values $\bar{f}(\mathbf{z}_{1}),\ldots,\bar{f}(\mathbf{z}_{Q})$. Since each candidate solution is feasible, its re-estimated objective values $\bar{f}(\cdot)$ provides an upper bound on $f^{*}$. Therefore, the approximated upper bound on the optimality gap for each candidate solution is $\bar{f}(\cdot)-\underaccent{\bar}{f}$. Finally, select the best solution $\mathbf{z}^{*}$ from the candidates, which is the one with the smallest optimality gap.
It is interesting to compare C-SAA with the approach proposed in Bertsimas and Kallus (2020). A key idea in Bertsimas and Kallus (2020) is to use the historial data (the dataset $\mathbb{D}$ in this paper, as the samples for the stochastic optimization and to weight them properly. Unfortunately, this is not possible in this problem $\mathcal{P}$ considered here, because only observational data is available, i.e., only the costs of the selected options are revealed. One way to proceed would be to define the sample set for the cost vector as all possible combinations of data points across the selection options. However, as the number of selection options and historical data points increase, this approach suffers from the curse of dimensionality and becomes intractable for large dataset with big selection space, which is the case in the application considered in this paper.
3.5 Contextual Robust Optimization (C-RO)
The performance of C-SAA largely hinges on the accuracy of the estimated conditional distribution. To mitigate the risk of model misspecification in forecasting, robust optimization can be employed to account for worst-case scenarios, making it a viable approach for contextual stochastic optimization. Let $\mathcal{U}(\mathbf{s}_{o})$ represent the contextual uncertainty set for the cost vector $\tilde{\mathbf{c}}_{o}$. The Contextual Robust Optimization (C-RO) problem for (3) is then formulated as follows:
(C-RO) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\max_{\bm{\xi}_{o}\in\mathcal{U}(\mathbf{s}_{o})}g(\mathbf{z}_{o},\bm{\xi}_{o}).$ (5)
To effectively leverage the conditional distributional forecasts and create tractable reformulations of Problem (5), two types of uncertainty sets are considered: a discrete uncertainty set $\mathcal{U}_{d}(\mathbf{s}_{o})$ and a budgeted interval uncertainty set $\mathcal{U}_{b}(\mathbf{s}_{o})$.
3.5.1 Contextual Discrete Uncertainty Set
With a discrete uncertainty set $\mathcal{U}_{d}(\mathbf{s}_{o})$, Problem (5) can be reformulated into a single-level problem by enumerating all possible values of $\tilde{\mathbf{c}}_{o}$ within $\mathcal{U}_{d}(\mathbf{s}_{o})$. This is a common technique in robust optimization models (Goerigk and Khosravi 2024):
[…]
s.t. $t\geq g(\mathbf{z}_{o},\bm{\xi}_{o}),\quad\forall\bm{\xi}_{o}\in\mathcal{U}_{d}(\mathbf{s}_{o}).$
A discrete distribution can be generated as mentioned before, by combining all the datapoints for the selection options. However, this will make the problem intractable as mentioned earlier. For comparison purposes, the results will provide a sampling counterpart to that approach in the same spirit as C-SAA. Since it is not quite satisfactory, another uncertainty set is considered next.
3.5.2 Contextual Budgeted Interval Uncertainty Set
A commonly used uncertainty set is the budgeted interval set $\mathcal{U}_{b}(\mathbf{s}_{o})$, which is defined by an uncertainty budget $B$. This budget limits the extent to which all components of the uncertain parameter $\tilde{\mathbf{c}}_{o}$ can simultaneously deviate to their upper bounds. This section assumes that the uncertain cost $\tilde{\mathbf{c}}_{o}$ is element-wise positive and $g(\mathbf{z}_{o},\tilde{\mathbf{c}}_{o})$ is a piecewise linear function of the following form:
$g(\mathbf{z}_{o},\tilde{\mathbf{c}}_{o})=(\tilde{\mathbf{c}}_{o}+\mathbf{a}_{o}^{j})^{T}\mathbf{z}_{o}\quad\text{if}~{}\mathbf{z}_{o}\in[\mathbf{b}^{j-1}_{o},\mathbf{b}^{j}_{o}],~{}j=1,\ldots n,$ (7)
where $\{\mathbf{b}^{0}_{o},\mathbf{b}^{1}_{o},\ldots,\mathbf{b}^{n}_{o}\}$ are the breakpoints of the piecewise linear function, and $\mathbf{a}_{o}^{j}=(a_{ivo}^{j})_{i\in\mathcal{I},v\in\mathcal{V}}$ is a deterministic cost vector corresponding to the segment $[\mathbf{b}^{j-1}_{o},\mathbf{b}^{j}_{o}]=[(b_{ivo}^{j-1})_{i\in\mathcal{I},v\in\mathcal{V}},(b_{ivo}^{j})_{i\in\mathcal{I},v\in\mathcal{V}}]$. Given these assumptions, the typical approach to ensure tractability when using a budgeted interval set is to dualize the inner maximization problem, resulting in the robust counterpart of the original problem (Bertsimas and Sim 2003, Bayram et al. 2022, Goerigk and Khosravi 2024).
Let the conditional prediction interval of $\tilde{\mathbf{c}}_{vo}$ be $[\bm{\underline{\xi}}_{vo},\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo}]$, where $\bm{\underline{\xi}}_{vo}=(\underline{\xi}_{ivo})_{i\in\mathcal{I}}$ and $\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo}=(\underline{\xi}_{ivo}+\hat{\xi}_{ivo})_{i\in\mathcal{I}}$ represent the lower and upper estimated conditional quantiles of the cost vector, respectively. The budgeted interval set $\mathcal{U}_{b}(\mathbf{s}_{o})$ can then be expressed as follows:
$\mathcal{U}_{b}(\mathbf{s}_{o},B)=\{\bm{\xi}_{o}\in\mathbb{R}^{|\mathcal{I}||\mathcal{V}|}:\xi_{ivo}=\underline{\xi}_{ivo}+\hat{\xi}_{ivo}\delta_{iv},\sum_{i\in\mathcal{I},v\in\mathcal{V}}\delta_{iv}\leq B,\delta_{iv}\in[0,1]\}.$ (8)
For a budgeted uncertainty set, it should be noted that learning the entire conditional distribution is unnecessary. Instead, learning the lower and upper conditional quantiles is sufficient to construct the conditional prediction interval. For instance, to create a 95% conditional prediction interval, only the 2.5% and 97.5% conditional quantiles are required. Let $q^{low}$ and $q^{upp}$ be the lower and upper quantiles for the conditional prediction interval of selection option $v$, respectively. Then, the lower bound $\bm{\underline{\xi}}_{vo}$ is defined as $\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid{\mathbf{s}_{vo}})\geq q^{low}\}$, and the upper bound $(\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo})$ is defined as $\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid{\mathbf{s}_{vo}})\geq q^{upp}\}$.
With a fixed $\mathbf{z}$, the inner maximization problem of Problem (5) is expressed as follows:
$\max\quad$ $\sum_{i\in\mathcal{I}}\sum_{v\in\mathcal{V}}\hat{\xi}_{ivo}z_{ivo}\delta_{iv}$ (9)
s.t. $\sum_{i\in\mathcal{I},v\in\mathcal{V}}\delta_{iv}\leq B$ (10)
$\delta_{iv}\in[0,1],\quad\forall i\in\mathcal{I},v\in\mathcal{V}.$ (11)
By taking the dual of the inner maximization problem and introducing auxiliary variables, a tractable reformulation of the robust counterpart of Problem (5) is obtained as follows:
[…]
s.t. $\pi+\lambda_{iv}\geq\hat{\xi}_{ivo}z_{ivo},\quad\forall i\in\mathcal{I},v\in\mathcal{V}$
$\pi\geq 0$
$\lambda_{iv}\geq 0,\quad\forall i\in\mathcal{I},v\in\mathcal{V}$
$z_{ivo}\leq b^{j}_{ivo}+M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$z_{ivo}\geq b^{j-1}_{ivo}-M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$w_{ivo}^{j}\geq(\underline{\xi}_{ivo}+a_{ivo}^{j})z_{ivo}-M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$y_{j}\in\{0,1\},\quad\forall j\in[n]$
$w_{ivo}^{j}\geq 0,\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
where $\pi$ and $(\lambda_{iv})_{i\in\mathcal{I},v\in\mathcal{V}}$ are the dual variables corresponding to Constraints (10) and (11), respectively, and $M$ is a reasonably large constant.
4 Application: The Ominichannel Multi-Courier Order Fulfillment Problem
This sections introduces a real-world order fulfillment problem encountered by the industrial partner. The proposed CSO framework is then applied to this problem, demonstrating significant advantages over the current practice and common baseline methods.

### v2

@@ 1 Introduction
Most existing studies on order fulfillment optimization have not explicitly considered the uncertainty in delivery timeliness. However, Freedman (2019) highlight the strong desire of customers for accurate delivery time promises and faster service. The survey indicates that around 44% of online shoppers have abandoned their shopping carts because the items they wanted would not arrive on time, while 20% abandon orders with unclear delivery dates. Furthermore, 28% of shoppers are willing to pay for expensive, expedited shipping to receive products at their desired time. The statistics clearly show that customers prioritize service quality, sometimes even more than the product value, when making purchase decisions. Moreover, Salari et al. (2022) point out the importance of considering delivery time uncertainty in order fulfillment to better meet customer expectations and improve satisfaction. Their proposed promised delivery time policy demonstrates a 6.1% increase in sales volume in numerical experiments. Notably, overpromising delivery times can lead to higher return rates, lower repurchase rates, and dissatisfaction with delayed service. Additionally, customers can have mixed sentiments over underpromised delivery times, as early deliveries may not align with their availability. Therefore, missing the exact promised delivery date can negatively impact short-term and long-term sales, both implicitly and explicitly (Salari et al. 2022, Cui et al. 2024). Given the crucial role of delivery timeliness, the paper incorporates delivery time uncertainty into the omnichannel multi-courier order fulfillment optimization problem through a contextual stochastic optimization framework.
One of the key challenges of Omnichannel Multi-Courier Order Fulfillment Optimization comes from the nature of historical data: only observational data—i.e., delivery time deviations for the fulfillment options actually chosen—are available. Counterfactual deviations for unchosen options are never observed, similar to the bandit feedback setting (Lattimore and Szepesvári 2020). The availability of observational data only, raises fundamental challenges that are typically not addressed in prior work.
@@ 1.1 Contributions
The key novelty in this paper is a CSO framework for an omnichannel multi-courier order fulfillment optimization with observational data only. Existing studies on CSO typically assume the availability of a dataset $\mathcal{D}=\{(\mathbf{s}_{1},\mathbf{c}_{1}),(\mathbf{s}_{2},\mathbf{c}_{2}),\ldots\}$, pairing covariates with fully observed uncertain parameter vectors. However, real-world scenarios often involve high-dimensional uncertain parameters where individual data points capture only partial information, particularly when uncertainty measurement depends on the chosen solution, which is the case in the application considered in this paper.
@@
    Generic CSO framework for observational data: The paper presents a novel, distribution‑agnostic contextual stochastic optimization (CSO) framework that converts problems with missing‑counterfactual observational data into tractable problems. A contextual distribution oracle is learned from partial observations, and the resulting problem is solved via two scalable solution paradigms—Contextual Sample Average Approximation (C‑SAA) and Contextual Robust Optimization (C‑RO).
@@
    Tailored contextual distribution learning for e-commerce delivery time deviations: The paper adapts calibrated probabilistic multi‐class classification and tree‐based quantile regression to exploit the discrete, ordinal nature of delivery time deviations, yielding well‑calibrated contextual distributions for the CSO framework.
@@
    Consolidation‑ and timeliness‑aware MILP for omnichannel, multi‑courier fulfillment: The paper proposes the first data‑driven mixed‑integer linear program (MILP) that jointly selects fulfillment centers and carriers and captures item‑consolidation discounts while embedding learned contextual deviation distributions in a tractable way.
@@
    Industry‐scale real‐world validation: The paper includes an extensive case study at a major e‐commerce retailer to (1) validate the efficacy of the proposed approaches in reducing expected fulfillment costs and (2) show significantly improvements in on‐time delivery rates compared to standard heuristics and baseline models. These results demonstrate the potential for substantial real‑world impact.
@@ 1.2 Structure of the Paper
The remainder of this paper is organized as follows. Section 2 reviews related literature. Section 3 formulates the omnichannel multi‑courier order‑fulfillment problem under delivery‑time deviation uncertainty. Section 4 develops a generic contextual stochastic optimization framework for observational data and introduces two tractable solution paradigms—Contextual Sample Average Approximation and Contextual Robust Optimization. Section 5 describes the machine‑learning methods (multi‑class classification and tree‑based quantile regression) used to implement the contextual distribution oracle. Section 6 presents the case study, computational results, and actionable insights from a real dataset. Finally, Section 7 concludes and outlines directions for future research.
@@ 2.1 Data-Driven Decision Making
This paper adds to the growing body of literature on data-driven decision-making, particularly at the intersection of machine learning and optimization, which has led to the development of CSO. Sadana et al. (2024) put forth a comprehensive survey on contextual optimization. Tutorials on various methodologies and applications of CSO models in logistics and operations management are available in the works of Mišić and Perakis (2020), Qi and Shen (2022), and Tian et al. (2023). Among the body of research on CSO, the most relevant to this paper are those that focus on learning contextual distributions of uncertain parameters and incorporating these forecasts into optimization models. Data-driven approaches typically build these forecasts from empirical residual errors (Deng and Sen 2022, Kannan et al. 2023, Perakis et al. 2023) or weighted empirical distributions based on proximity to the training data (Bertsimas and Kallus 2020, Notz and Pibernik 2022). However, these studies assume full observation of uncertain parameters for each data point, which differs from the context of this paper. Recent research has also explored integrating downstream optimization effects into ML model training. Elmachtoub and Grigas (2022) introduce the Smart Predict then Optimize (SPO) loss and its convex surrogate for point prediction models, with subsequent extensions to contextual distributional estimation (Qi et al. 2021, Kallus and Mao 2023). However, the SPO loss framework is not directly applicable to this paper’s context, as the dataset analyzed here is observational. In such datasets, outcomes are only partially observed based on decisions made, presenting unique challenges not addressed by these existing methods.
This paper thus contributes to the field of data-driven decision-making that handles observational data, a domain explored by prior research, e.g., in Bertsimas et al. (2019) and Jo et al. (2021), using optimal prescriptive trees for prescriptive analytics in personalized treatment. While these methods effectively learn policies mapping features to actions, they cannot be directly applied to the fulfillment problem considered in this paper due to the presence of instance-specific hard constraints on decisions. The proposed framework addresses these limitations, offering a novel approach to CSO problems with observational data and constrained decision spaces.
@@ 2.2 Order Fulfillment Problem
3 The Omnichannel Multi-Courier Order Fulfillment Problem
This section introduces a real-world order fulfillment problem encountered by the industrial partner.